OpenStudy (epicteatime):

If a meterstick moves to your right at a constant velocity of 0.8c, what would you measure its length to be?

OpenStudy (epicteatime):

@nincompoop

OpenStudy (epicteatime):

anyoneee

OpenStudy (epicteatime):

@aaronq

OpenStudy (astrophysics):

Length contraction

OpenStudy (astrophysics):

\[\huge L = \frac{ L_0 }{ \gamma } = L_0 \sqrt{1-\frac{ v^2 }{ c^2 }}\]

OpenStudy (astrophysics):

\[L_0\] is your proper length, L is the length observed by an observer in relative motion (respect to the object), v is your relative velocity, and finally c is the speed of light

OpenStudy (epicteatime):

yep that xD

OpenStudy (epicteatime):

I think the answers 0.6m, what do you think?

OpenStudy (astrophysics):

\[L_0 = 1m\] \[v = 0.8c\] \[L = 1m \times \sqrt{1-\frac{ (0.8c)^2 }{ c^2 }} = \sqrt{1-\frac{ 0.64c^2 }{ c^2 }}m = 0.6m ~~ \checkmark\]

OpenStudy (epicteatime):

HUDDAH!

OpenStudy (epicteatime):

xD Thanks for le confirmation fwendo :D

OpenStudy (astrophysics):

Np :)

OpenStudy (epicteatime):

but just to be sure, do you cancel out the c^2 or smt?

OpenStudy (astrophysics):

Yes, it gets cancelled out

OpenStudy (epicteatime):

ah, thanks xD

OpenStudy (astrophysics):